# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 1.1 Model and main result
where if $\alpha=1$ and $l(\infty)=\infty$, then it can be chosen that ${\cal R}(t)=l(t)^{-(\frac{1}{2K}-\varepsilon)}$; if $\alpha=1$ and $l(\infty)<\infty$, then it can be chosen that ${\cal R}(t)=\ell(t)^{-(\frac{1}{2}-\varepsilon)}$; and if $\alpha<1$, then it can be chosen that ${\cal R}(t)=t^{-\eta}$.
To establish Theorem 1.1, we require two key ingredients as in (1.2). One is an analysis of the density function $\rho_{1}(t)$ (see Theorem 1.2), and the other is an asymptotic result for the non-collision probability of of $N$ independent heavy-tailed random walks with fixed starting point ${\bf x}$ (see Theorem 1.4).
Theorem 1.2 (Density asymptotics)
Under the assumptions of Theorem 1.1, the density of CRW at time $t$ is $\rho_{1}(t):=\rho_{1}(0,t)={\mathbb{E}}[\xi(0)]$. Then, there exists some $\eta>0$ such that for any $\varepsilon>0$,
$\rho_{1}(t)=\frac{g_{\alpha}(0)\hskip 1.0ptl(2t)}{t}\big{(}1+O\big{(}{\cal R}(t)\big{)}\big{)},$ (1.10)
where if $\alpha=1$ and $l(\infty)=\infty$, then it can be chosen that ${\cal R}(t)=l(t)^{-(\frac{1}{2K}-\varepsilon)}$; if $\alpha=1$ and $l(\infty)<\infty$, then it can be chosen that ${\cal R}(t)=\ell(t)^{-1/2}$; and if $\alpha<1$, then it can be chosen that ${\cal R}(t)=t^{-\eta}$.
Remark 1.3
The most interesting case is the recurrent underlying random walk with $\alpha=1$, $l(\infty)=\infty$ and $B_{t}\in{\cal B}$ in Theorems 1.1 and 1.2. This mirrors the 2-dimensional simple CRW studied in [BG80, CP04, LTZ18]. Indeed, in the 2D setting with our notation one takes $l(t)=\ell(t)=\log t$ and $K=1$, and then $\rho_{1}(t)$ and $\rho_{N}(t)$ exhibit the same leading asymptotics as in Theorems 1.1 and 1.2. Moreover, the error in $\rho_{1}(t)$ can be improved from $O\big{(}l(t)^{-(1/2-\varepsilon)}\big{)}$ with any $\varepsilon>0$ to $O\big{(}l(t)^{-1/2}\big{)}$.
There exists a natural coupling between coalescing random walks and independent random walks so that the two collections of paths are identical up to the first collision time. For convenience, the following theorem is stated for $N$ independent random walks.
Theorem 1.4 (Non-collision probability asymptotics)
Let ${\bf x}\in{\mathbb{Z}}^{N}$, and let $({\mathbf{X}}_{t})_{t\geq 0}=(X^{1}_{t},\ldots,X^{N}_{t})_{t\geq 0}$ be $N$ independent continuous-time random walks with jump rate $1$ and symmetric jump kernel $p(\cdot)$ as in (1.4), starting from ${\mathbf{X}}_{0}={\bf x}$. Define the first collision time
$\tau_{{\bf x}}:=\inf\{t\geq 0:X^{i}_{t}=X^{j}_{t}~{}\text{for some}~{}1\leq i<j\leq N\},$ (1.11)
and set the non-collision probability $p_{\rm NC}(t):={\mathbb{P}}(\tau_{{\bf x}}>t)$. Assume that $B_{t}\in{\cal B}$. Then, there exists $c_{\bf x}$ only depending on ${\bf x}$ such that
$p_{\rm NC}(t)=\frac{c_{\bf x}}{l(2t)^{\upsilon_{N}}}\Big{(}1+O({\cal R}(t))\Big{)},$ (1.12)
where if $\alpha=1$, then it can be chosen that ${\cal R}(t)=\int_{t}^{\infty}\frac{1}{u\hskip 1.0pt\ell(u)^{2-\varepsilon}}\mathrm{d}u=o\big{(}\ell(t)^{-2+\varepsilon}\big{)}$ with any $\varepsilon>0$; and if $10<\alpha<$ 1, then it can be chosen that ${\cal R}(t)=t^{-\eta}$ with some $\eta>1/\alpha-1$.
1.2 Proof strategy
We first outline the proof of Theorems 1.2. For high-dimensional ($d\geq 3$) coalescing random walks, van der Berg and Kesten [vdBK00, vdBK02] developed a framework to analyze the asymptotics of the system density $\rho_{1}(t)$ by deriving an ODE through the generator calculations. This robust method has since been applied for CRW on transitive graphs [HLYZ22] and to coalescing systems with multiple particle types [TZ24]. Lukins, Tribe, and Zaboronski [LTZ18] adapted this framework to the barely recurrent case in two dimensions. Furthermore, it is adapted by Lukins, Tribe and Zaboronski [LTZ18] to two-dimensional coalescing random walks which is barely recurrent. In our setup, we start by applying the generator (1.3) to the density function to obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=-\sum_{y}p(y){\mathbb{P}}(\xi_{t}(0,y)=11)$ (1.13)
To have particles both at sites 00 and $y$ at time $t$, we can trace back the trajectories for their ancestors (which are typically unique) over a time interval $s$ (with $s\ll t$). In effect, we consider these as two backward random walks that have not coalesced between $t-s$ and $t$. By applying the Markov property at time $t-s$, we can decompose
${\mathbb{P}}(\xi_{t}(0,y)=11)\displaystyle\approx\sum_{z_{1},z_{2}}{\mathbb{P}}(\xi_{t-s}(z_{1},z_{2})=11)\hskip 1.0pt{\mathbb{P}}(\widehat{X}^{0,y}_{s}=(z_{1},z_{2}),\hat{\tau}_{y}>s),$ (1.14)
where $(\widehat{X}^{0,y}_{t})_{t\geq 0}$ denotes two independent backward random walks starting from 00 and $y$, and $\hat{\tau}_{y}$ is their first collision time. Choosing $s=s(t)$ with $1\ll s\ll t$ appropriately ensures that the backward walks are typically far apart at time $s$ (i.e., $|z_{1}-z_{2}|\asymp B_{s(t)}$), so that ${\mathbb{P}}(\xi_{t-s}(z_{1},z_{2})=11)\approx{\mathbb{P}}(\xi_{t-s}(z_{1})=1)\times{\mathbb{P}}(\xi_{t-s}(z_{2})=1)=\rho_{1}^{2}(t-s)$. Substituting this into (1.14) and using $\rho_{1}(t-s)\approx\rho_{1}(t)$ (since $t-s$ is close to $t$) yields
${\mathbb{P}}(\xi_{t}(0,y)=11)\approx\rho_{1}^{2}(t-s){\mathbb{P}}(\hat{\tau}_{y}>s)\approx\rho_{1}^{2}(t){\mathbb{P}}(\hat{\tau}_{y}>t),$ (1.15)
where we used the fact that the random walk is either transient or barely recurrent (in the sense that $B_{t}\in{\cal B}$) so that ${\mathbb{P}}(\hat{\tau}_{y}>s)\approx{\mathbb{P}}(\hat{\tau}_{y}>t)$ for a suitable choice of $s=s(t)$. Combining (1.13) and (1.15) produces a simple ODE for $\rho_{1}(t)$.
Note that (1.14)-(1.15) shows an approximation for the two-point function. One can extend these ideas to general ${\bf x}\in{\mathbb{Z}}^{N}$ to prove Theorem 1.1. First, we can establish the asymptotic independence
${\mathbb{P}}(\widehat{X}^{{\bf x}}_{s}={\bf z},\hat{\tau}_{{\bf x}}>s)\approx{\mathbb{P}}(\widehat{X}^{{\bf x}}_{s}={\bf z}){\mathbb{P}}(\hat{\tau}_{{\bf x}}>s),$ (1.16)
$\rho_{N}({\bf x},t)\approx\sum_{\bf z}\rho_{N}({\bf z},t-s){\mathbb{P}}(\widehat{X}^{{\bf x}}_{s}={\bf z},\hat{\tau}_{{\bf x}}>s)\approx\rho_{1}^{N}(t-s)\hskip 1.0pt{\mathbb{P}}(\hat{\tau}_{{\bf x}}>s)\approx\rho_{1}^{N}(t)\hskip 1.0pt{\mathbb{P}}(\hat{\tau}_{{\bf x}}>t),$ (1.17)
It remains to establish Theorem 1.4, which concerns the non-collision probability $p_{\rm NC}(t)={\mathbb{P}}(\tau_{{\bf x}}>t)$ for $N$ independent random walks $({\mathbf{X}}_{t})_{t\geq 0}$. The first step is to derive a crude bound Lemma 3.2, so that we can choose a suitable $s=s(t)\ll t$ satisfying $p_{\rm NC}(s)\approx p_{\rm NC}(t)$. Then, we derive an ODE for $p_{\rm NC}(t)$ in a manner analogous to the previous arguments, as also done in [LTZ18] for 2D simple CRW. Since the decrease in $p_{\rm NC}(t)$ is due to collisions between any two random walks $X^{i}$ and $X^{j}$, by the symmetry of $p$ and translation invariance we have
$-\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)&=&\displaystyle\sum_{i<j}\sum_{x_{i},x_{j}}2p(x_{i}-x_{j}){\mathbb{P}}_{{\bf x}}(\tau_{{\bf x}}>t,X^{i}_{t}=x_{i},X^{j}_{t}=x_{j})\\[10.0pt] &=&\displaystyle\sum_{i<j}\sum_{y_{ij}}2p(y_{ij}){\mathbb{P}}_{{\bf x}}(\tau_{{\bf x}}>t,\Delta^{ij}_{t}=y_{ij}),\end{array}$ (1.18)
@@ 1.2 Proof strategy
where $\Delta^{ij}_{t}:=X^{i}_{t}-X^{j}_{t}$ and $y_{ij}:=y_{i}-y_{j}$. Decompose at time $s$ and evolve $\Delta^{ij}$ backward from $t$ to $s$. As only $\Delta^{ij}_{t}$ is fixed, other free pairs of walks are unlikely to collide form $s$ to $t$. Thus,
${\mathbb{P}}_{{\bf x}}(\tau_{{\bf x}}>t,\Delta^{ij}_{t}=y_{ij})\approx\sum_{z_{ij}}{\mathbb{P}}_{{\bf x}}(\Delta^{ij}_{s}=z_{ij},\tau_{{\bf x}}>s){\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{t-s}=z_{ij},\hat{\tau}_{ij}>t-s).$ (1.19)
@@
Applying asymptotic independence to ${\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{t-s}=z_{ij},\hat{\tau}_{ij}>t-s)$ and noting $|y_{ij}|,|z_{ij}|\ll t^{1/\alpha}$, the second term in the r.h.s. of (1.19) can be replaced by ${\mathbb{P}}_{0}(\widehat{\Delta}^{ij}_{t-s}=0)\hskip 1.0pt{\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s)$. This, together with (1.18) and a priori estimate $p_{\rm NC}(s)\approx p_{\rm NC}(t)$, implies that
@@
$-\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)\displaystyle\approx p_{\rm NC}(t)\cdot\sum_{i<j}\sum_{y_{ij}}2p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s)\hskip 1.0pt{\mathbb{P}}_{0}(\widehat{\Delta}^{ij}_{t-s}=0).$ (1.20)
@@ 1.3 Discussion
To the best of our knowledge, Theorem 1.1 is the first result to investigate the $N$-point function for coalescing heavy-tailed random walks, revealing a rich structure of spatial correlations in these systems. Along the way, we establish several sharp estimates for heavy-tailed random walks (particularly in the regime $\alpha\leq 1$), which have not appeared in the literature before. We derive quantitative error bounds in asymptotics for key quantities such as the transition probability, the potential kernel, and first hitting times. These estimates are of independent interest and may serve as valuable tools for future investigations into heavy-tailed processes.
Looking ahead, one promising direction is to study the $N$-point function for $\alpha>1$. In this regime, the approach of deriving an ODE becomes inapplicable due to the recurrence of the underlying random walk, and new methods must be developed. Another direction is to investigate the scaling limit of the system. By rescaling space by $\varepsilon$ and time by $\varepsilon^{\alpha}$ and letting $\varepsilon\downarrow 0$, the resulting particle system initially starts from every point of ${\mathbb{R}}$. It is known [EMS13] that if $\alpha>1$, the rescaled system exhibits the phenomenon of coming down from infinity, namely the particles become locally finite at any positive time. However, this does not occur in our setting. Instead, one may consider the particles as a closed subset of ${\mathbb{R}}$. We conjecture that, in a suitable topology, the limiting set (as $\varepsilon\downarrow 0$) at a positive time has a fractional Hausdorff dimension of $1-\alpha$, as suggested by Theorem 1.1.
@@ 2 Heavy-tailed random walks
Let $(X^{x}_{t})_{t\geq 0}$ be the heavy-tailed continuous-time random walk starting from $X^{x}_{0}=x$. Throughout this section, we assume that the jump rate is 1 and that the symmetric jump kernel $p(\cdot)$ has the form (1.4). For brevity, when $x=0$ we also write $(X^{0}_{t})_{t\geq 0}$ as $(X_{t})_{t\geq 0}$. By [IL71, Thm 2.6.5] and its proof, the characteristic function of $p$ is given by
$\phi(u)=\exp\big{\{}-\gamma|u|^{\alpha}\widetilde{L}\big{(}|u|^{-1}\big{)}\big{\}},$ (2.1)
@@
$\phi_{t}(u)={\mathbb{E}}\big{[}{\mathbb{E}}[e^{iuX_{t}}|N_{t}]\big{]}=\exp\big{(}-t(1-\phi(u))\big{)}.$ (2.2)
We will repeatedly use the following two results for a SVF $H(t)$. The first is the Potter bound [BGT89, Thm 1.5.6]: There exists $c>0$ such that for any $t\geq s\geq c$ and any $\varepsilon>0$,
$(t/s)^{-\varepsilon}\leq H(t)/H(s)\leq(t/s)^{\varepsilon}.$ (2.3)
The second result is Karamata’s theorem [BGT89, Prop 1.5.8-1.5.10]: For any $c\geq 0$
$\begin{array}[]{ll}\text{For}~{}\alpha\in(0,1),&\displaystyle\int_{t}^{\infty}\frac{H(s)}{s^{1/\alpha}}\mathrm{d}s\sim\frac{1/\alpha-1}{t^{1/\alpha-1}}H(t);\\[10.0pt] \text{For}~{}\alpha=1,&\displaystyle\int_{t}^{\infty}\frac{H(s)}{s}\mathrm{d}s~{}\text{is a SVF if it is finite, and~{}}\frac{1}{H(t)}\int_{t}^{\infty}\frac{H(s)}{s}\mathrm{d}s\to 0;\\[10.0pt] &\displaystyle\int_{c}^{t}\frac{H(s)}{s}\mathrm{d}s~{}\text{is a SVF, and~{}}\frac{1}{H(t)}\int_{c}^{t}\frac{H(s)}{s}\mathrm{d}s\to 0;\\[10.0pt] \text{For}~{}\alpha\in(1,2),&\displaystyle\int_{c}^{t}\frac{H(s)}{s^{1/\alpha}}\mathrm{d}s\sim\frac{t^{1-1/\alpha}}{1-1/\alpha}H(t).\end{array}$ (2.4)
To avoid potential ambiguities, below for $k$-th power of a function $f(t)$, we write $f(t)^{k}$ instead of $f^{k}(t)$. For example, in Section 4 we do not write $\rho_{1}^{N-1}(t-s)$ but $\rho_{1}(t-s)^{N-1}$.
@@ 2.1 Transition probability
One key input to the ODE framework is a sharp asymptotic estimate of the transition probability ${\mathbb{P}}(X_{t}=x)$. While Gnedenko’s LLT provides an optimal $o(1)$ error in approximating the transition probability by that of the limiting $\alpha$-stable process, our analysis requires a finer quantitative error bound. This motivates the following proposition on the difference between transition probabilities at distinct spatial points.
@@
$\big{|}{\mathbb{P}}(X_{t}=x+y)-{\mathbb{P}}(X_{t}=x)\big{|}\leq\frac{C|y|}{B_{t}^{2}}.$ (2.5)
@@
$\begin{array}[]{l}\displaystyle\big{|}{\mathbb{P}}(X_{t}=x+y)-{\mathbb{P}}(X_{t}=x)\big{|}\leq\frac{|y|}{2\pi B_{t}^{2}}\int_{-\pi B_{t}}^{\pi B_{t}}|u\cdot\phi_{t}(u/B_{t})|\mathrm{d}u.\end{array}$ (2.6)
Recalling $\phi_{t}(u)=\exp\{-t(1-\phi(u))\}$ from (2.2) with $\phi(u)$ given in (2.1), we have
$|\phi_{t}(u/B_{t})|=\exp\Big{\{}-t\Big{(}1-e^{-\frac{\gamma|u|^{\alpha}\widetilde{L}(B_{t}/u)}{B_{t}^{\alpha}}}\Big{)}\Big{\}}.$ (2.7)
By the Potter bound (2.3), for any $0<\varepsilon<\alpha$,
$\frac{\gamma|u|^{\alpha}}{t}\frac{t\widetilde{L}(B_{t})}{B_{t}^{\alpha}}\frac{\widetilde{L}(B_{t}/u)}{\widetilde{L}(B_{t})}\geq C\frac{\gamma|u|^{\alpha}}{t}\frac{1}{u^{\varepsilon}}=\frac{C\gamma u^{\alpha-\varepsilon}}{t},$
@@
Since $|u|e^{-C\gamma|u|^{\alpha-\varepsilon}}$ is integrable over ${\mathbb{R}}$, the bound (2.6) implies the desired result.
@@
$\phi^{\text{ns}}(u)=\exp\big{\{}icu-\gamma|u|^{\alpha}\widetilde{L}\big{(}|u|^{-1}\big{)}\big{(}1-i\beta\text{sgn}(u)\omega(u,\alpha)\big{)}\big{\}}=\phi(u)e^{i\hskip 1.0pt\widetilde{\omega}(u)}.$ (2.9)
@@
Large deviation estimates. For $\alpha\in(0,2)$, we recall the following large deviation and local large deviation estimates from [Ber19] (which were originally established for discrete-time random walks, but can be adapted to continuous-time via standard Poissonian arguments):
[…]
@@
Let $s\ll t$. For the random walk $(X^{x}_{t})_{t\geq 0}$, there exists a constant such that for any $\varepsilon>0$,
$\big{|}{\mathbb{P}}(X^{x}_{t}=0)-{\mathbb{P}}(X^{x}_{t+s}=0)\big{|}\leq\frac{C}{B_{t}\hskip 1.0pt(t/s)^{\frac{1}{1+\alpha}-\varepsilon}}.$ (2.11)
Proof Using $1=\sum_{z}{\mathbb{P}}(X_{s}^{0}=z)=\sum_{z}{\mathbb{P}}(X_{s}^{z}=0)$, we rewrite
${\mathbb{P}}(X_{t}^{x}=0)-{\mathbb{P}}(X^{x}_{t+s}=0)=\sum_{z}\big{(}{\mathbb{P}}(X_{t}^{x}=0)-{\mathbb{P}}(X_{t}^{x}=z)\big{)}{\mathbb{P}}(X_{s}^{z}=0).$
Define $u:=t/s\gg 1$. We discuss in two regions: $|z|\leq B_{t}u^{-\kappa}$ and $|z|>B_{t}u^{-\kappa}$ for some $\kappa<1/\alpha$. For any $\varepsilon>0$, since $B_{t}u^{-\kappa}/B_{s}\geq u^{-\kappa}(t/s)^{1/\alpha-\varepsilon}\to\infty$, the large deviation implies
$\sum_{|z|>B_{t}u^{-\kappa}}{\mathbb{P}}(X_{s}^{z}=0)\leq\frac{CsL(B_{t}u^{-\kappa})}{B_{t}^{\alpha}u^{-\alpha\kappa}}=\frac{tL(B_{t})}{B_{t}^{\alpha}}\hskip 1.0pt\frac{CL(B_{t}^{\alpha}u^{-\kappa})}{u^{1-\alpha\kappa}L(B_{t})}\leq\frac{C}{u^{1-\alpha\kappa-\varepsilon}},$
where we used the Potter bound in the last inequality. Therefore,
$\sum_{|z|>B_{t}u^{-\kappa}}\big{|}{\mathbb{P}}(X_{t}^{x}=0)-{\mathbb{P}}(X_{t}^{x}=z)\big{|}{\mathbb{P}}(X_{s}^{z}=0)\leq\frac{C}{B_{t}}\sum_{|z|>B_{t}u^{-\kappa}}{\mathbb{P}}(X_{s}^{z}=0)\leq\frac{C}{B_{t}u^{1-\alpha\kappa-\varepsilon}}.$ (2.12)
In the other region, by Proposition 2.1,
$\sum_{|z|\leq B_{t}u^{-\kappa}}\big{|}{\mathbb{P}}(X_{t}^{x}=0)-{\mathbb{P}}(X_{t}^{x}=z)\big{|}{\mathbb{P}}(X_{s}^{z}=0)\leq\frac{CB_{t}u^{-\kappa}}{B_{t}^{2}}\sum_{z}{\mathbb{P}}(X_{s}^{z}=0)\leq\frac{C}{B_{t}u^{\kappa}}.$ (2.13)
Taking $\kappa=1/(1+\alpha)$, the desired result follows from the bounds (2.13) and (2.12).
For non-symmetric kernel $p$, if we replace $X$ by suitably centered random walk, then the large deviation remains valid [Ber19] and thus (2.11) also holds.

### v2

@@ 1.2 Proof strategy
where $\Delta^{ij}_{t}:=X^{i}_{t}-X^{j}_{t}$ and $y_{ij}:=y_{i}-y_{j}$. Decompose at time $s$ and evolve $\Delta^{ij}$ backward from $t$ to $s$. As only $\Delta^{ij}_{t}$ is fixed, other free pairs of walks are unlikely to collide from $s$ to $t$. Thus,
${\mathbb{P}}_{{\bf x}}(\tau_{{\bf x}}>t,\Delta^{ij}_{t}=y_{ij})\approx\sum_{z_{ij}}{\mathbb{P}}_{{\bf x}}(\Delta^{ij}_{s}=z_{ij},\tau_{{\bf x}}>s){\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{t-s}=z_{ij},\hat{\tau}_{ij}>t-s).$ (20)
@@
Applying asymptotic independence to ${\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{t-s}=z_{ij},\hat{\tau}_{ij}>t-s)$ and noting $|y_{ij}|,|z_{ij}|\ll t^{1/\alpha}$, the second term in the r.h.s. of (20) can be replaced by ${\mathbb{P}}_{0}(\widehat{\Delta}^{ij}_{t-s}=0)\hskip 1.0pt{\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s)$. This, together with (19) and a priori estimate $p_{\rm NC}(s)\approx p_{\rm NC}(t)$, implies that
@@
$\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)\displaystyle\approx p_{\rm NC}(t)\cdot\sum_{i<j}\sum_{y_{ij}}2p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>t-s)\hskip 1.0pt{\mathbb{P}}_{0}(\widehat{\Delta}^{ij}_{t-s}=0).$ (21)
@@ 1.3 Discussion
To the best of our knowledge, Theorem 1.3 is the first result to investigate the $N$-point function for coalescing heavy-tailed random walks, revealing a rich structure of spatial correlations in these systems. Along the way, we establish several sharp estimates for heavy-tailed random walks (particularly in the regime $\alpha\leq 1$), which have not appeared in the literature before. We derive quantitative error bounds in asymptotics for key quantities such as the transition probability, the potential kernel, and first hitting times. These estimates are of independent interest and may serve as valuable tools for future investigations into heavy-tailed processes. Together, these contributions constitute the main novelty of the present work.
In the marginally recurrent case, the assumption that $B_{t}\in\mathcal{B}$ is of a technical nature. More precisely, to obtain a quantitative estimate of the non-collision probability $p_{\rm NC}(t)$ in Theorem 1.2, an intermediate bound of the type established in Lemma 3.1 is required, where one of the error terms is expressed in terms of $\ell(t)^{-\kappa}$. If the condition $B_{t}\in\mathcal{B}$ fails so that $\ell(t)^{K}\leq l(t)$ for all $K$, then in (64) the “error” term would, in contrast, dominate the “main” term $p_{\rm NC}(t)$ which only involves a power of $l(t)$. Nevertheless, we expect that the main result should remain valid for $\alpha=1$ even when $B_{t}\notin\mathcal{B}$. But in that case the proof would require an a priori crude bound on $p_{\rm NC}(t)$ in terms of $l(t)$, which would in turn demand new ideas and techniques. We therefore leave this as an open problem for future work. We also note a possible extension to the case of non-symmetric jump kernels $p(\cdot)$. As discussed in Remark 2.8, the present proof strategy should still apply provided that Proposition 2.4 can be established for the potential kernel in the non-symmetric setting.
Looking ahead, one promising direction is to study the $N$-point function for $\alpha>1$. In this regime, the approach of deriving an ODE becomes inapplicable due to the recurrence of the underlying random walk. One possible avenue worth exploring is to establish a Karlin–McGregor formula for heavy-tailed random walks and to perform an asymptotic analysis of this formula, as is done for coalescing Brownian motions in MRTZ (06). However, because heavy-tailed paths exhibit crossings caused by jumps, new methods would be required to derive an appropriate Karlin–McGregor formula in this setting. Another direction is to investigate the scaling limit of the system. By rescaling space by $\varepsilon$ and time by $\varepsilon^{\alpha}$ and letting $\varepsilon\downarrow 0$, the resulting particle system initially starts from every point of $\mathbb{R}$}. It is known EMS (13) that if $\alpha>1$, the rescaled system exhibits the phenomenon of coming down from infinity, namely the particles become locally finite at any positive time. However, this does not occur in our setting where $0<\alpha\leq 1$. Instead, one may consider the particles as a closed subset of $\mathbb{R}$}. We conjecture that, in a suitable topology, the limiting set (as $\varepsilon\downarrow 0$) at a positive time has a fractional Hausdorff dimension of $1-\alpha$, as suggested by Theorem 1.3.
@@ 2 Heavy-tailed random walks
Let $(X^{x}_{t})_{t\geq 0}$ be the heavy-tailed continuous-time random walk starting from $X^{x}_{0}=x$. Throughout this section, we assume that the jump rate is 1 and that the symmetric jump kernel $p(\cdot)$ has the form (4). For brevity, when $x=0$ we also write $(X^{0}_{t})_{t\geq 0}$ as $(X_{t})_{t\geq 0}$. By (IL, 71, Thm 2.6.5) and its proof, the characteristic function of $p$ is given by
$\phi(u)=\exp\big\{-\gamma|u|^{\alpha}\widetilde{L}\big(|u|^{-1}\big)\big\},$ (22)
@@
$\phi_{t}(u)={\mathbb{E}}\big[{\mathbb{E}}[e^{iuX_{t}}|N_{t}]\big]=\exp\big(-t(1-\phi(u))\big).$ (23)
We will repeatedly use the following two results for a SVF $H(t)$. The first is the Potter bound (BGT, 89, Thm 1.5.6): There exists $c>0$ such that for any $t\geq s\geq c$ and any $\varepsilon>0$,
$(t/s)^{-\varepsilon}\leq H(t)/H(s)\leq(t/s)^{\varepsilon}.$ (24)
The second result is Karamata’s theorem (BGT, 89, Prop 1.5.8-1.5.10): For any $c>0$,
[…]
@@ 2.1 Transition probability
One key input to the ODE framework is a sharp asymptotic estimate of the transition probability ${\mathbb{P}}(X_{t}=x)$. While Gnedenko’s LLT provides an optimal $o(1)$ error in approximating the transition probability by that of the limiting $\alpha$-stable process, our analysis requires a finer quantitative error bound. This motivates the following proposition on the difference between transition probabilities at distinct spatial points. As a supplementary remark, we note that refined local limit theorems with quantitative error bounds are available when the random walk is assumed to lie in the domain of normal attraction of an $\alpha$-stable distribution (i.e., the SVF $\widetilde{L}$ in (22) is constant and $B_{t}=t^{1/\alpha}$), see CJR (23) and references therein.
@@
$\big|{\mathbb{P}}(X_{t}=x+y)-{\mathbb{P}}(X_{t}=x)\big|\leq\frac{C|y|}{B_{t}^{2}}.$ (26)
@@
$\begin{array}[]{l}\displaystyle\big|{\mathbb{P}}(X_{t}=x+y)-{\mathbb{P}}(X_{t}=x)\big|\leq\frac{|y|}{2\pi B_{t}^{2}}\int_{-\pi B_{t}}^{\pi B_{t}}|u\cdot\phi_{t}(u/B_{t})|\mathrm{d}u.\end{array}$ (27)
Recalling $\phi_{t}(u)=\exp\{-t(1-\phi(u))\}$ from (23) with $\phi(u)$ given in (22), we have
$|\phi_{t}(u/B_{t})|=\exp\Big\{-t\Big(1-e^{-\frac{\gamma|u|^{\alpha}\widetilde{L}(B_{t}/|u|)}{B_{t}^{\alpha}}}\Big)\Big\}\leq\exp\Big\{-\frac{c\gamma|u|^{\alpha}t\widetilde{L}(B_{t}/|u|)}{B_{t}^{\alpha}}\Big\},$ (28)
where in the inequality we used that for small $x>0$, $1-e^{-x}\geq cx$ for some $c>0$. By (6) and the Potter bound (24), for any $0<\varepsilon<\alpha$,
$\frac{t\widetilde{L}(B_{t})}{B_{t}^{\alpha}}\frac{\widetilde{L}(B_{t}/|u|)}{\widetilde{L}(B_{t})}\geq C|u|^{-\varepsilon},$
@@
Since $|u|e^{-C\gamma|u|^{\alpha-\varepsilon}}$ is integrable over $\mathbb{R}$}, the bound (27) implies the desired result.
@@
$\phi^{\text{ns}}(u)=\exp\big\{icu-\gamma|u|^{\alpha}\widetilde{L}\big(|u|^{-1}\big)\big(1-i\beta\text{sgn}(u)\omega(u,\alpha)\big)\big\}=\phi(u)e^{i\hskip 1.0pt\widetilde{\omega}(u)}.$ (30)
@@
Large deviation estimates. For $\alpha\in(0,2)$, we recall the following large deviation and local large deviation estimates from Ber (19) (which were originally established for discrete-time random walks, but can be adapted to continuous-time via standard Poissonian arguments):
$\begin{array}[]{cl}\text{Large deviation}&\quad{\mathbb{P}}(|X^{0}_{t}|>y)\leq\frac{Ct\hskip 1.0ptL(y)}{(1+y)^{\alpha}}\quad{\rm as}~y/B_{t}\to\infty.\\[8.0pt] \text{Local large deviation}&\quad{\mathbb{P}}(X^{0}_{t}=y)\leq\frac{Ct\hskip 1.0ptL(|y|)}{B_{t}(1+|y|)^{\alpha}}\quad{\rm for~all}~y\in{\mathbb{Z}}.\end{array}$ (31)
@@
Let $s\ll t$. For the random walk $(X^{x}_{t})_{t\geq 0}$, there exists a constant such that
$\big|{\mathbb{P}}(X^{x}_{t}=0)-{\mathbb{P}}(X^{x}_{t+s}=0)\big|\leq\frac{C}{B_{t}}\hskip 1.0pt(s/t).$ (32)
Proof It suffices to show that
$\big|\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(X^{x}_{t}=0)\big|\leq\frac{C}{tB_{t}},$
since the desired result then follows by integrating this inequality from $t$ to $t+s$. Observe that
$\frac{\mathrm{d}}{\mathrm{d}t}\phi_{t}(u)=\frac{\mathrm{d}}{\mathrm{d}t}e^{-t(1-\phi(u))}=-(1-\phi(u))\hskip 1.0pte^{-t(1-\phi(u))},$
Applying the inverse Fourier transform yields
$\Big|\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(X^{x}_{t}=0)\Big|=\Big|\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-iux}\frac{\mathrm{d}}{\mathrm{d}t}\phi_{t}(u)\mathrm{d}u\Big|\leq\frac{1}{2\pi}\int_{-\pi}^{\pi}(1-\phi(u))\hskip 1.0pte^{-t(1-\phi(u))}\mathrm{d}u.$
Using the inequality $xe^{-x}\leq e^{-x/2}$ for all $x\geq 0$, and setting $x=t(1-\phi(u))$, we obtain
$\Big|\frac{\mathrm{d}}{\mathrm{d}t}{\mathbb{P}}(X^{x}_{t}=0)\Big|\leq t^{-1}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-\frac{t}{2}(1-\phi(u))}\mathrm{d}u=t^{-1}{\mathbb{P}}(X^{x}_{t/2}=0),$
where the last equality follows from the identity $e^{-\frac{t}{2}(1-\phi(u))}=\phi_{t/2}(u)$ and the inverse Fourier transform formula. This concludes the proof by the local limit theorem.
For non-symmetric kernel $p$, if we replace $X$ by suitably centered random walk, then the large deviation remains valid Ber (19) and thus (32) also holds.